OpenStudy (anonymous):

What is the amplitude, period, and phase shift of f(x) = -4 sin(2x + π) - 5? (2 points)

OpenStudy (anonymous):

amplitude = -4; period = 2π; phase shift: x = -PI/2 amplitude = -4; period = π; phase shift: x = PI/2 amplitude = 4; period = π; phase shift: x = -PI/2 amplitude = 4; period = 2π; phase shift: x = PI/2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):